OpenStudy (anonymous):

What do I do next? Problem below! It's law of sines. I always give out metals and become a fan!

OpenStudy (anonymous):

\[\frac{ \sin 130 }{ d}=\frac{ \sin 20 }{ 35 }\]

OpenStudy (anonymous):

sin20 * d = 35 sin 130 d= [\frac{ 35 \sin 130 }{ \sin 20 }\] plug it into ur calculator

OpenStudy (anonymous):

d = 3.08

OpenStudy (anonymous):

i got 15.6?

OpenStudy (anonymous):

are you on radian or degree

OpenStudy (anonymous):

degrees, but when I switch to radians I still don't get that. Anyways neither answers are right because my diagram shows it should be bigger than 35

OpenStudy (anonymous):

i just did it agian on my calculator..and I get 78.39

OpenStudy (anonymous):

That seems like the right answer! HOw did you do it?

OpenStudy (anonymous):

i plugged 35sin130 and i get the answer and just divide the answer by sin20 have to put the ")" after20

OpenStudy (anonymous):

I got! Thanks so much!!! :)

OpenStudy (anonymous):

no problem